Good Evening:) here is my question:
The roots of a quadratic equation are -1 and 3. What is its defining equation?
Can you also Give the solution?

Answer by srinivas.g(540)

give roots = -1 , 3
sum of the roots = -1 +3
=2
product of the roots = -1*3
= -3
quadratic equation :
formula : x^2 - sum of the roots (x) + product of the roots = 0
x^2 -2x+ (-3) =0
x^2-2x-3 =0